OpenStudy (anonymous):

When –3 is multiplied by the sum of a number and 8, the result is –21. Find the number.

OpenStudy (anonymous):

-3(x + 8) = -21 -- distribute through the parenthesis -3x - 24 = -21 -- add 24 to each side -3x = -21 + 24 -3x = 3 -- divide by -3 x = -1

OpenStudy (anonymous):

any questions ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Meg opened a savings account with $44.50. She puts in $22.25 in the account each week when she gets paid. She is saving to purchase a blu-ray player that costs $178.00. How many weeks will it take her to save enough money? A. 6 weeks B. 7 weeks C. 8 weeks D. 9 weeks

OpenStudy (anonymous):

44.5 + 22.25w = 178 can you do this ? Just solve for w , the number of weeks.

OpenStudy (anonymous):

are you having problems ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thats 8 weeks??

OpenStudy (anonymous):

hold on..let me check

OpenStudy (anonymous):

lets sub your answer in and see if it is true 44.5 + 22.25(8) = 178 44.5 + 178 = 178 that is not correct 44.5 + 22.25w = 178 subtract 44.5 from both sides to get 22.25w by itself what do you get from that ?

OpenStudy (anonymous):

9 week then

OpenStudy (anonymous):

44.5 - 44.5 + 22.25w = 178 - 44.5 you see how the 44.5's will cancel each other out on the left side now combine like terms 22.25w = 133.5 do you understand so far ?

OpenStudy (anonymous):

22.25w = 133.5 now divide both sides by 22.25 to get w by itself 22.25w/22.25 = 133.5/22.25 now what do you get ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

lol thanks

OpenStudy (anonymous):

i got another question

OpenStudy (anonymous):

check it.. 44.5 + 22.25(6) = 178 44.5 + 133.50 = 178 178 = 178 (correct) answer is 6 weeks

OpenStudy (anonymous):

can u help me out?

OpenStudy (anonymous):

possibly...what you got ?

OpenStudy (anonymous):

When modeling word problems as mathematical equations, which of the following is a synonym for the division operation? A. difference B. decreased C. per D. combined

OpenStudy (anonymous):

i think its C but im not sure

OpenStudy (anonymous):

synonym means the same, right ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

well..not sure...none of them make sense I know it is not decreased or difference....combined means addition....so I guess it is per.

OpenStudy (anonymous):

difference is subtraction, decreased is subtraction, combined is addition, and "percent" means "per 100" and represents dividing by 100

OpenStudy (anonymous):

your right....it is per

OpenStudy (anonymous):

you need anymore help ?

OpenStudy (anonymous):

im fine thanks